OpenStudy (anonymous):

Write the comparison as a ratio in simplest form, using a fraction. 6 to 12

OpenStudy (anonymous):

1:2

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

6/12 = 3/6 == 1/2 = 1:2

OpenStudy (anonymous):

== (is the same as)

OpenStudy (anonymous):

6/12 = 3/6 = 1/2 >>>>> 1:2

OpenStudy (anonymous):

thank you